How long will it take a 16 g sample of bismuth-210 to decay such that 2.0 g of the element remains?
1
Expert's answer
2012-10-01T10:10:13-0400
Half life of bismuth-210 is 5 days.

m=m(0)*exp(-b*t)

m/m(0) = 2 g
/ 16 g = 1/8
1/8 = (1/2)^3 = 0.125
=> 3 half life periods

5
days x 3 = 15 days